# Controllable Motion Generation via Diffusion Modal Coupling

### v1

@@ IV-A Modal Coupling
()($\downarrow$)( ↓ )
meanFDE
()($\downarrow$)( ↓ )
minMR
()($\downarrow$)( ↓ )
meanMR
()($\downarrow$)( ↓ )
minOR
()($\downarrow$)( ↓ )
meanOR
()($\downarrow$)( ↓ )
VBD, $S=1$ 1.707 1.707 4.593 4.593 0.551 0.551 0.170 0.170
VBD, $S=3$ 1.265 1.717 3.294 4.597 0.394 0.549 0.126 0.172
VBD, $S=9$ 0.968 1.708 2.422 4.557 0.279 0.544 0.102 0.178
cVBD, $k=9$, $S=1$ 1.425 1.425 3.667 3.667 0.437 0.437 0.173 0.173
cVBD, $k=9$, $S=3$ 1.025 1.418 2.530 3.650 0.290 0.430 0.126 0.178
Ours, $S=1$ $k=3$,   $\delta=8$ 1.267 1.267 3.523 3.523 0.378 0.378 0.107 0.107
$k=3$,   $\delta=15$ 1.216 1.216 3.221 3.221 0.367 0.367 0.117 0.117
$k=9$,   $\delta=8$ 1.019 1.019 2.678 2.678 0.260 0.260 0.070 0.070
$k=9$,   $\delta=15$ 0.940 0.940 2.423 2.423 0.212 0.212 0.057 0.057
Ours, $S=3$ $k=3$, $\delta=8$ 1.093 1.281 2.986 3.565 0.299 0.382 0.090 0.109
$k=3$, $\delta=15$ 1.117 1.219 2.937 3.223 0.322 0.364 0.103 0.139
$k=9$, $\delta=8$ 0.892 1.017 2.323 2.662 0.210 0.262 0.050 0.065
$k=9$, $\delta=15$ 0.857 0.938 2.190 2.419 0.182 0.209 0.046 0.060
TABLE I: Quantitative validation for motion prediction accuracy on Waymo. We evaluate the proposed method using two configurations for the number of modes: $k=3$, which considers only steering modes, and $k=9$, which incorporates both steering and speed modes. $S$ denotes the number of sampled trajectories per scenario for each setting. Due to limited computing resources, the training size was reduced to one-tenth of its original size. We anticipate that the reported metrics would improve further if trained on the full dataset.
IV-B Prior Parametrization
Next, we present our parametrization of the proposed multi-modal prior distribution. As shown in Figure 3, it is crucial to keep the prior modes well-separated (i.e., not overlapping) so that the diffusion model can clearly distinguish the target mode from the others. Additionally, each mode must be constrained from generating excessively large samples that could complicate the diffusion process.
To specify the Gaussian mean of each prior mode, we draw on the concept of placing evenly spaced points on a high-dimensional sphere. Let $d$ be the data dimension, and $k$ the number of modes. If $k\leq d+1$, which is often the case in complex data distributions suited to diffusion models, this arrangement can be achieved by placing the vertices of a $(k-1)$-simplex in $\mathbb{R}^{d}$ [14]. Let $e_{1:k}\subset\mathbb{R}^{k}$ be the standard basis vectors. We define the means $\mu_{1:k}$ as follows:
$w_{i}$ $=\delta\cdot\sqrt{\frac{k}{k+1}}\cdot(e_{i}-\mathbbm{1}_{k}\cdot\frac{1}{k})$ (17)
$\mu_{i}$ $=[w_{i}^{1},...,w_{i}^{k},0,...,0]^{T}\in\mathbb{R}^{d},$ (18)
where $\delta$ is a hyper-parameter for adjusting the sphere radius and $w_{i}^{j}$ denotes the $j$-th component of $w_{i}$. One can verify that $\|\mu_{i}\|^{2}=\delta$ for all $i\in[1,k]$, and that these points lie at a pairwise distance of $\delta\cdot\sqrt{2+2/(k-1)}$. Finally, we choose the variance of each Gaussian so that their confidence intervals are at a certain level do not overlap with respect to the above pairwise distance. That is, for all $i\in[1,k]$,
$\sigma_{i}\sqrt{\mathcal{X}^{2}_{d,c}}\leq\delta\cdot\sqrt{2+\frac{2}{(k-1)}},$ (19)
where $c$ denotes the desired confidence level and $\mathcal{X}^{2}_{d,c}$ is the $c$ quantile of the chi‐square distribution with $d$ degress of freedom  [15].
V Experiments
@@ V-A Waymo - Controllable Motion Prediction
We evaluate the proposed method on the motion prediction task using the Waymo Open Motion Dataset [16] that features diverse and realistic autonomous driving scenarios. Because motion prediction is inherently probabilistic and multi-modal, it is crucial to capture an unbiased distribution of possible future trajectories. In our experiments, we focus on the single-agent setting where distinct modes in the target distribution are more clearly defined, thus better highlighting the advantages of our approach.
We adopt Versatile Behavior Diffusion (VBD) [12] as the backbone of our model. A query-centric, Transformer-based scene encoder first processes the scene context, including scenario information such as traffic lights, map polylines, and agent trajectories, to capture interrelationships among all scene components. Building on this contextual understanding, a denoiser model then predicts a clean control sequence from a noised input. Under the single-agent setting, the denoiser considers only the ego-agent’s control trajectory, while other agents remain visible to the scene encoder, given their potential influence on the ego-agent’s future behavior.
Next, we describe the parameter settings of the proposed method. We adopt a two-layered mode design: the first layer corresponds to steering modes (left turn, right turn, and go straight), while the second layer addresses speed control (acceleration, deceleration, and maintaining speed). If both layers are considered, we take the Cartesian product of these two layers, which yields $k=9$ total modes. During training, we set $\sigma_{i}=1$ for all $i\in[1,k]$. The data dimension is $80$, corresponding to $2$D control predictions over $40$ future steps (i.e., $4$ seconds).
@@ V-A1 Motion Prediction Accuracy
Table I summarizes the primary metrics used to evaluate prediction accuracy on the test set. We report four standard metrics. Average Displacement Error (ADE) measures the mean distance between the ground-truth future ego-agent trajectories and the model predictions, averaged over the future time horizon. Final Displacement Error (FDE) measures the displacement error at the final time step. Miss Rate (MR) measures the recall of the trajectory predictions. And OffRoad (OR) assesses prediction consistency in how often the predicted trajectories drive off the road. For each metric, we report the minimum and mean values computed from the sampled trajectories. In addition, we implement a variant of VBD, referred to as conditioned VBD (cVBD), which is given explicit mode labels as part of its input.
Although enhancing prediction performance is not the primary goal of this project, Table I demonstrates notable improvements in the reported metrics, especially in the mean statistics. We attribute these gains primarily to our use of a multi-modal prior distribution, mirroring the positive results observed on the 2D toy set (Figure 3). This benefit is further underscored by the performance increase when the number of modes $k$ rises from $3$ to $9$. Meanwhile, the configurations with $\delta=15$ outperform those with $\delta=8$, emphasizing the importance of ensuring that prior mode distributions do not overlap. Lastly, as cVBD still relies on a unimodal prior distribution, its ability to effectively learn each mode in the target distribution is limited. By contrast, our method leverages strong modal coupling to better capture multi-modal behavior, ultimately leading to more robust predictions.
Method
minADE
()($\uparrow$)( ↑ )
meanADE
()($\uparrow$)( ↑ )
minOR
()($\downarrow$)( ↓ )
meanOR
()($\downarrow$)( ↓ )
ACC[ST]
()($\uparrow$)( ↑ )
ACC[SP]
()($\uparrow$)( ↑ )
ACC
()($\uparrow$)( ↑ )
VBD+G $\omega=1$ 1.049 1.814 0.070 0.108 0.877 0.450 0.387
$\omega=10$ 1.790 2.717 0.070 0.113 0.952 0.624 0.601
$\omega=100$ 2.294 3.443 0.121 0.184 0.944 0.711 0.701
cVBD,     $k=9$ 1.996 2.901 0.111 0.205 0.883 0.806 0.707
Ours $k=9$, $\delta=8$ 1.932 2.233 0.035 0.050 0.906 0.702 0.647
$k=9$, $\delta=15$ 2.162 2.369 0.023 0.026 0.917 0.701 0.662
TABLE II: Quantitative validation for trajectory synthesis controllability, using $9$ sampled trajectories per scenario. The test set consists of $140$ scenarios, each manually labeled with potentially feasible ego-agent futures that deviate from the dataset’s original trajectories. Note that higher ADE values imply a greater deviation from the reference trajectories, not indicating the degraded prediction performance.
Figure 4: Qualitative results for motion prediction. In standard diffusion, trajectories are sampled randomly from a unimodal prior, offering no inherent controllability. cVBD takes the intended mode as part of its model input, but still employs a unimodal prior, preventing it from effectively presenting distinct data modes. VBD-G applies guidance to steer generation, but it relies heavily on the guidance influence factor, making it difficult to balance sample fidelity against controllability. Our method integrates modal coupling with a multi-modal prior distribution, yielding notable improvements in both sample fidelity and controllability.
@@ V-A2 Controllable Trajectory Synthesis
We assess controllability in Table II by measuring how well sampled trajectories remain realistic and feasible despite deviating from the reference trajectories (Figure 4). To generate Table II, we manually label potential ego-agent futures with an emphasis on steering feasibility. For instance, if the ego-agent is in a lane that allows turning, yet the reference trajectory continues straight, we count the turning maneuver as feasible. However, identifying valid speed modes that differ from the reference is more challenging, and thus, occasional mislabeling may occur. The final evaluation set comprises $140$ scenarios, which we provide in our codebase   [^1] https://github.com/RobinWangSD/Diffusion-with-Multi-Modal-Priors.git.
In addition to the conditioned VBD baseline, we examine a post-hoc guidance strategy for the unconditioned VBD model, referred to as VBD+G. At each denoising step $t\in[2,T]$, we update the posterior mean:
$\hat{\mu}_{\theta}(x_{t})$ $=\mu_{\theta}(x_{t})+\omega\cdot\mathop{\nabla}\limits_{\mu_{\theta}(x_{t})}f(x^{0}_{\theta}\Big{(}\mu_{\theta}(x_{t}),t-1\Big{)}),$ (20)
where $\mu_{\theta}(x_{t})$ is the posterior mean estimate computed from the predicted $x_{\theta}^{0}(x_{t},t)$ via (6), $\omega$ is a parameter controlling guidance strength, and $f$ is the guidance function applied to the predicted $x_{0}$, treating $\mu_{\theta}(x_{t})$ as $x_{t-1}$. These steps follow the procedures described in [12]. In our experiments, $f$ aligns with the mode list used during training for both the proposed method and the conditioned VBD. Specifically, given a target steer-speed label pair, $f$ penalizes any input trajectory that exceeds the specified steer-speed range.
To evaluate performance, we continue using Average Displacement Error (ADE) with respect to reference trajectories, which measures how far sampled trajectories differ from the reference. We also include OffRoad (OR), an important metric that captures fidelity and feasibility of sampled trajectories. Finally, we introduce an overall accuracy (ACC) metric to measure how often the samples match the intended steer-speed mode, and further decompose it into separate measures for steering (ACC[ST]) and speed (ACC[SP]).
As shown in Table II, the guidance-based method struggles to balance sample fidelity and controllability. Although increasing $\omega$ does improve overall accuracy, it also significantly worsens OffRoad. In contrast, our proposed method consistently maintains high sample fidelity while achieving reasonably high accuracy. Furthermore, as $\delta$ increases, thus separating each prior mode more distinctly, performance continues to improve. It is worth noting that VBD-G requires extensive backpropagation through the diffusion model, making its generation noticeably slower than our method which simply performs a standard reverse process from the specified prior mode. Meanwhile, the conditioned VBD achieves comparable performance to VBD-G at $\omega=100$.
@@ V-B Maze2d - Multi-task Control
The proposed method naturally extends to multi-task control. Rather than confining control-level constraints to a single task, we can also view each task itself as a separate mode. This perspective allows our method to compose multiple tasks within a single, unconditioned diffusion model. We illustrate this with evaluations in Maze2D, where a unified model is trained to perform long-horizon path planning across various maze configurations.
@@
We define the diffusion model to predict $384$-step state-control trajectories conditioned on the given initial and goal positions. Given that the state and control dimensions are $3$ and $2$, respectively, the target distribution has dimension $1920$. We set $\sigma_{i}=1$ for $i\in[1,3]$, and choose $\delta=30$. As shown in Table III, the proposed method achieves performance comparable to the baseline across all tested maze configurations. Note that our aim is not to outperform the existing methods. Instead, we demonstrate that by treating each distinct task as an individual data mode, the proposed method establishes effective modal coupling over a multi-modal prior, which makes it possible to handle various tasks using only a single unconditioned diffusion model.
@@ VI-B Proof of Lemma 2
From (35) to (36), the constant terms that do not involve $x_{t-1}$ are all omitted. Following the standard Gaussian density function, the mean and variance of $p(x_{t-1}|x_{t},x_{0})$ can be parameterized as $\mathcal{N}(\tilde{\mu},\tilde{\beta})$ where:
@@
$+\frac{\mu}{(1-\bar{\alpha}_{t})\eta_{T}}\cdot\frac{-(1-\bar{\alpha}_{t})}{\sqrt{\alpha_{t}}}$
$=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x_{t}-\frac{\mu}{\eta_{T}}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\sigma\epsilon_{t}\Big{)}$ (44)

### v2

@@ V Experiments
$(\downarrow)$
Collision
$(\downarrow)$
OffRoad
$(\downarrow)$
ACC[ST]
$(\uparrow)$
ACC[SP]
$(\uparrow)$
ACC
$(\uparrow)$
Inference Time
$(\downarrow)$
CG $\eta=1$ 2.614 5.999 0.054 0.127 0.882 0.968 0.746 2.717
$\eta=10$ 2.862 6.655 0.057 0.160 0.916 0.987 0.902 2.703
$\eta=100$ 4.681 12.358 0.059 0.275 0.976 0.996 0.970 2.712
CFG $\lambda=0.9$ 3.994 10.194 0.070 0.162 0.878 0.833 0.746 1.535
$\lambda=1$ 3.034 7.138 0.069 0.159 0.877 0.977 0.864 1.566
$\lambda=1.05$ 2.687 6.055 0.065 0.157 0.885 0.996 0.883 1.544
$\lambda=1.1$ 3.099 7.506 0.062 0.184 0.887 1.000 0.887 1.547
Ours $\delta=0.5$ 2.506 5.104 0.053 0.121 0.895 0.999 0.894 0.899
$\delta=1.0$ 2.266 4.757 0.043 0.092 0.914 1.000 0.914 0.892
$\delta=2.0$ 2.525 5.303 0.489 0.101 0.907 0.998 0.904 0.906
$\delta=4.0$ 2.772 5.697 0.060 0.135 0.895 0.993 0.888 0.864
EM 2.578 5.500 0.060 0.140 0.897 0.995 0.893 0.891
TABLE I: Quantitative results on the Waymo Open Motion Dataset under reference driving modes. Driving modes comprise steering and speed categories. minADE/minFDE are averaged across steering–speed modes, mitigating the effect of imbalanced mode frequencies.
Figure 4: Qualitative results for motion prediction. In standard diffusion, trajectories are sampled randomly from a unimodal prior, offering no inherent controllability. CG applies guidance to steer generation, but it relies heavily on the guidance influence factor, making it difficult to balance sample fidelity against controllability. CFG blends unconditional and conditional outputs for guidance, which limits controllability and fidelity when the target lies off the reference data manifold. Our method integrates modal coupling with a multi-modal prior distribution, yielding notable improvements in both sample fidelity and controllability.
Method
Collision
$(\downarrow)$
OffRoad
$(\uparrow)$
ACC
$(\uparrow)$
CG $\eta=1$ 0.050 0.125 0.595
$\eta=10$ 0.060 0.135 0.830
$\eta=100$ 0.060 0.205 0.960
CFG $\lambda=0.9$ 0.080 0.215 0.725
$\lambda=1$ 0.075 0.195 0.875
$\lambda=1.05$ 0.060 0.185 0.900
$\lambda=1.1$ 0.060 0.195 0.910
Ours $\delta=0.5$ 0.060 0.115 0.915
$\delta=1.0$ 0.040 0.110 0.930
$\delta=2.0$ 0.055 0.120 0.920
$\delta=4.0$ 0.070 0.140 0.925
EM 0.060 0.180 0.905
TABLE II: Quantitative validation for trajectory synthesis controllability, using 88 sampled trajectories per scenario. The test set consists of $200$ scenarios, each manually labeled with potentially feasible ego-agent futures that deviate from the dataset’s original trajectories.
We begin with the Waymo dataset, showing that our approach produces realistic and feasible motions while preserving controllability, in contrast to post-hoc guidance baselines. We then show that our method extends naturally to a multi-task setting by treating each task as a distinct mode. Unlike prior diffusion-based planners that train a separate model per task [22, 23], a single model suffices for our approach while maintaining strong per-task performance.
@@ V-A Waymo - Controllable Motion Prediction
We evaluate on the Waymo Open Motion Dataset [24], under single-agent setting with explicitly-defined mode structure. Following [25], ground-truth actions $L$ are deterministically labeled along two independent axes: steering (left, right, straight) and speed (accelerate, decelerate, maintain speed) with a differentiable classifier $L=g(x)$. The task is to generate an $8$-s ego control trajectory consistent with the queried action label while respecting map topology and contextual agent interactions. We predict $2$-D controls over $80$ future steps (dimension $160$), yielding $k=9$ action modes from the Cartesian product of the two axes.
Versatile Behavior Diffusion (VBD) [12] is adopted as the model backbone. From VBD, we reuse: (i) the query-centric Transformer encoder for scene-level interaction encoding, and (ii) diffusion timesteps embeddings for in-context conditioning. Our instantiation departs from VBD in two ways: (1) we flatten the control sequence as a single vector representation [16], and (2) during denoising, we apply cross-attention from the ego control-sequence embeddings to the scene context to capture ego-environment interactions.
We compare against two guidance baselines. Classifier Guidance (CG) applies the deterministic classifiers (used for labeling) to yield gradients with respect to the unconditioned score network [14]. Classifier-Free Guidance (CFG) trains additional conditional models with shared parameters [19]. Both methods apply guidance in a post-hoc manner, with standard formulations given as follows:
$\mathbf{CG}\quad$ $\hat{\epsilon}_{\theta}(x_{t},t)=\epsilon_{\theta}(x_{t},t)+\eta\nabla_{x_{t}}g(\mu_{\theta}(x_{t})),$
$\mathbf{CFG}\quad$ $\hat{\epsilon}_{\theta}(x_{t},t)=\epsilon_{\theta}(x_{t},t)+\lambda(\epsilon_{\theta}(x_{t},t,L)-\epsilon_{\theta}(x_{t},t)),$
Unless stated otherwise, we set per-mode scales $\sigma_{i}=1$ for $i\in\{1,\dots,k\}$. We also evaluate an empirical-means (EM) initialization that sets the multi-modal prior’s means and standard deviations from training-set statistics.
Next, we describe the evaluation metrics used in our study. We report Average Displacement Error (ADE), i.e., the mean Euclidean distance between predicted and ground-truth positions, and Final Displacement Error (FDE), i.e., the terminal position error. To assess physical plausibility, we compute Collision and Off-Road rates[26], defined as the fraction of rollouts that intersect another agent or leave the drivable area, respectively. We introduce accuracy (ACC) statistics to evaluate control fidelity, while further decomposing it into per-axis statistics: ACC[ST] (steering) and ACC[SP] (speed).
@@ V-A1 Motion Prediction Accuracy
Table I summarizes the prediction accuracy results, evaluated against ground-truth labels of the ego-agent’s future behaviors. While the focus of this work is on controllability, we demonstrate that coupling diffusion model training on explicitly labeled datasets with a multi-modal prior distribution substantially improves predictive performance.
Comparison to guidance baselines. For both guidance approaches (CG, CFG), the guidance coefficient ($\eta$ for CG; $\lambda$ for CFG) is sensitive to tuning and enforces a control-fidelity trade-off. Since these methods operate with a unimodal prior, they offer limited mode-specific learning and coverage. Even at their best settings, our method outperforms them: a mode-coupled prior allocates capacity per mode and enables direct, label-aligned control, yielding more robust predictions. Computationally, CG requires per-step backpropagation through the diffusion model, while CFG doubles forward passes. Both introduce higher inference cost than our single reverse process from the selected prior component.
Ablation on $\delta$ and EM prior. Setting $\delta=1.0$ consistently outperforms $\delta=0.5$, improving inter-component separation in the prior and reducing overlap among reverse-diffusion trajectories. Excessively large $\delta$ degrades performance: as in the $2$-D toy example (Figure 3), components displaced far from the origin yield implausible trajectory distributions. Under the EM configuration, steering means concentrate near zero because left/right-turn trajectories contain extended straight segments. Speed statistics are less skewed but still exhibit partial mode collapse. Although EM preserves cross-mode action-sequence correlations, the increased component overlap reduces separability and harms accuracy.
@@ V-A2 Controllable Trajectory Synthesis
Table II evaluates realism and feasibility when generation is conditioned on driving modes that deviate from the dataset reference. We manually annotate feasible future modes: e.g., if the ego is in a lane that permits turning while the reference continue straight, the turn is marked feasible. The evaluation set contains $200$ scenarios and will be released with the code. Visualizations of generated trajectories using various controllability methods on manually labeled scenarios are presented in Figure 4.
Under out-of-distribution conditioning, our method continues to produce feasible, on-road trajectories, as reflected in the low collision and off-road rates. CG/CFG often fail to realize off-reference modes with realistic generations: e.g., intersection scenes frequently go off-road after crossing. While achieving superior sample realism, the proposed method also maintains stable, high accuracy. These results underscore the limitations of guidance-based control (particularly post-hoc guidance) that our approach is designed to overcome.
@@ V-B Maze2d - Multi-task Control
We extend our method method to multi-task control. Rather than confining control-level constraints to a single task, we can also view each task itself as a distinct data mode. We illustrate this with evaluations in Maze2D, where a single, task-agnostic model is trained to perform long-horizon path planning across three maze configurations.
@@
We define the diffusion model to predict $384$-step state-control trajectories conditioned on the initial and goal positions. With $3$-D states and $2$-D controls, the target distribution is $1920$-dimensional. We set $\sigma_{i}=1$ for $i\in[1,3]$, and choose $\delta=30$. As shown in Table III, the unified model matches baseline performance across all tested mazes. Note that the goal here is not to surpass the baseline, but to demonstrate that casting tasks as modes establishes modal coupling over a multi-modal prior, enabling diverse tasks to be handled by a single model without task-specific conditioning achitectures.
@@ VI-B Proof of Lemma 2
[…]
@@
$\hskip-48.36958pt=\frac{1}{\sqrt{\alpha_{t}}}\Big(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\sigma\epsilon_{t}\Big)$ (44)
$\hskip-19.91684pt+\frac{\mu}{(1-\bar{\alpha}_{t})\eta_{T}}\cdot\frac{-(1-\bar{\alpha}_{t})}{\sqrt{\alpha_{t}}}$
$\hskip-48.36958pt=\frac{1}{\sqrt{\alpha_{t}}}\Big(x_{t}-\frac{\mu}{\eta_{T}}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\sigma\epsilon_{t}\Big)$ (45)